# Hallucination Detection in LLMs Using Spectral Features of Attention Maps

### v1

@@ 6 Ablation studies
To better understand the behaviour of our method under different conditions, we conduct a comprehensive ablation study. This analysis provides valuable insights into the factors driving the $\operatorname{LapEigvals}$ performance and highlights the robustness of our approach across various scenarios. In order to ensure reliable results, we perform all studies on the TriviaQA dataset, which has a reasonable input size and number of examples.
@@ 6.1 How does the number of eigenvalues influence performance?
First, we verify how the number of eigenvalues influences the performance of the hallucination probe and present results in Figure 4. Generally, using more eigenvalues improves performance, but there is less variation in performance among different values of $k$ for $\operatorname{LapEigvals}$. Moreover, $\operatorname{LapEigvals}$ achieves significantly better performance with smaller input sizes, as $\operatorname{AttnEigvals}$ with the largest $k=100$ fails to surpass $\operatorname{LapEigvals}$’s performance at $k=5$. These results confirm that spectral features derived from the Laplacian carry a robust signal indicating the presence of hallucinations and highlight the strength of our method.
Figure 4: Model performance across different values of top-$k$ eigenvalues: $k\in\{5,10,25,50,100\}$ for TriviQA dataset with $temp=1.0$.
@@ 6.2 Does using all layers at once improve performance?
Second, we demonstrate that using all layers of an LLM instead of a single one improves performance. In Figure 5, we compare per-layer to all-layer efficacy. For the per-layer approach, better performance is generally achieved in later LLM layers. Notably, peak performance varies across LLMs, requiring an additional search for each new LLM. In contrast, the all-layer probes consistently outperform the best per-layer probes across all LLMs. This finding suggests that information indicating hallucinations is spread across many layers of LLM, and considering them in isolation limits detection accuracy. Further, Table 4 in Appendix D summarises outcomes for the two variants on all datasets and LLM configurations examined in this work.
Figure 5: Analysis of model performance across different layers and LLMs for TriviQA dataset with $temp=1.0$ and $k=100$ top eigenvalues (results for models operating on all layers provided for reference).
@@ 6.3 Does sampling temperature influence results?
Here, we compare $\operatorname{LapEigvals}$ to baselines on hallucination datasets produced with several temperatures used during decoding. Higher temperatures typically produce more hallucinated examples (Lee, 2023; Renze, 2024), leading to dataset imbalance. Thus, to mitigate the effect of data imbalance, we sample a subset of $1000$ hallucinated and $1000$ non-hallucinated examples $10$ times for each temperature and train hallucination probes. Interestingly, in Figure 6, we observe that all models improve their performance at higher temperatures, but $\operatorname{LapEigvals}$ consistently achieves the best accuracy on all considered temperature values. The correlation of efficacy with temperature may be attributed to differences in the characteristics of hallucinations at higher temperatures compared to lower ones (Renze, 2024). Also, hallucination detection might be facilitated at higher temperatures due to underlying properties of softmax function (Veličković et al., 2024), and further exploration of this direction is left for future work.
Figure 6: Test AUROC for different sampling $temp$ values during answer decoding on the TriviaQA dataset, using $k=100$ eigenvalues for $\operatorname{LapEigvals}$ and $\operatorname{AttnEigvals}$ with the Llama-3.1-8B LLM. Error bars indicate the standard deviation over 10 balanced samples containing $N=1000$ examples per class.
@@ 6.4 How does $\operatorname{LapEigvals}$ generalizes?
To check whether our method generalises across datasets, we trained the hallucination probe on features from the training split of one QA dataset and evaluated it on the features from the test split of a different QA dataset. Due to space limitations, we present results for selected datasets and provide extended results and absolute efficacy values in Appendix E. Figure 7 showcases the percentage drop in Test AUROC when using a different training dataset compared to training and testing on the same QA dataset. We can observe that $\operatorname{LapEigvals}$ provides a performance drop comparable to other baselines, and in several cases, it generalises best. Interestingly, all methods exhibit poor generalisation on TruthfulQA, possibly due to dataset size or imbalance. Additionally, in Appendix E, we show that $\operatorname{LapEigvals}$ achieves the highest test performance in all scenarios (except for TruthfulQA).
Figure 7: Generalisation across datasets measured as a per cent performance drop in Test AUROC (less is better) when trained on one dataset and tested on the other. Training datasets are indicated in the plot titles, while test datasets are shown on the $x$-axis. Results computed on Llama-3.1-8B with $k=100$ top eigenvalues and $temp=1.0$. Results for all datasets are presented in Appendix E.
@@ 6.5 How does performance vary across prompts?
Lastly, to assess the stability of our method across different prompts used for answer generation, we compared the results of the hallucination probes trained on features from four distinct prompts, the content of which is included in Appendix F. As shown in Table 2, $\operatorname{LapEigvals}$ consistently outperforms all baselines across all four prompts. While we can observe variations in performance across prompts, $\operatorname{LapEigvals}$ demonstrates the lowest standard deviation ($0.05$) compared to $\operatorname{AttnLogDet}$ ($0.016$) and $\operatorname{AttnEigvals}$ ($0.07$), indicating its greater robustness.
Table 2: Test AUROC across four different prompts for answers on the TriviaQA dataset using Llama-3.1-8B  with $temp=1.0$ and $k=50$ (some prompts have led to fewer than 100 tokens). Prompt $\boldsymbol{p_{3}}$ was the main one used to compare our method to baselines, as presented in Tables 1.
@@ 7 Related Work
White-box methods have emerged as a promising approach for detecting hallucinations (Farquhar et al., 2024; Azaria and Mitchell, 2023; Arteaga et al., 2024; Orgad et al., 2025). These methods are universal across all LLMs and do not require additional domain adaptation compared to black-box ones (Farquhar et al., 2024). They draw inspiration from seminal works on analysing the internal states of simple neural networks (Alain and Bengio, 2016), which introduced linear classifier probes – models operating on the internal states of neural networks. Linear probes have been widely applied to the internal states of LLMs, e.g., for detecting hallucinations.
@@
One of the first such probes was SAPLMA (Azaria and Mitchell, 2023), which demonstrated that one could predict the correctness of generated text straight from LLM’s hidden states. Further, the INSIDE method (Chen et al., 2024) tackled hallucination detection by sampling multiple responses from an LLM and evaluating consistency between their hidden states using a normalised sum of the eigenvalues from their covariance matrix. Also, (Farquhar et al., 2024) proposed a complementary probabilistic approach, employing entropy to quantify the model’s intrinsic uncertainty. Their method involves generating multiple responses, clustering them by semantic similarity, and calculating Semantic Entropy using an appropriate estimator. To address concerns regarding the validity of LLM probes, (Marks and Tegmark, 2024) introduced a high-quality QA dataset with simple true/false answers and causally demonstrated that the truthfulness of such statements is linearly represented in LLMs, which supports the use of probes for short texts.
@@
Self-consistency methods (Liang et al., 2024), like INSIDE or Semantic Entropy, require multiple runs of an LLM for each input example, which substantially lowers their applicability. Motivated by this limitation, (Kossen et al., 2024) proposed to use Semantic Entropy Probe, which is a small model trained to predict expensive Semantic Entropy (Farquhar et al., 2024) from LLM’s hidden states. Notably, (Orgad et al., 2025) explored how LLMs encode information about truthfulness and hallucinations. First, they revealed that truthfulness is concentrated in specific tokens. Second, they found that probing classifiers on LLM representations do not generalise well across datasets, especially across datasets requiring different skills. Lastly, they showed that the probes could select the correct answer from multiple generated answers with reasonable accuracy, which they concluded with the LLM making mistakes at the decoding stage besides knowing the correct answer.
@@
Recent studies have started to explore hallucination detection exclusively from attention maps. (Chuang et al., 2024a) introduced the lookback ratio, which measures how much attention LLMs allocate to relevant input parts when answering questions based on the provided context. The work most closely related to ours is (Sriramanan et al., 2024), which introduces the $\operatorname{AttentionScore}$ method. Although the process is unsupervised and computationally efficient, the authors note that its performance can depend highly on the specific layer from which the score is extracted. We also demonstrate that it performs poorly on the datasets we evaluated. Nonetheless, we drew inspiration from their approach, particularly using the lower triangular structure of matrices when constructing features for the hallucination probe.
@@ 8 Conclusions
In this work, we demonstrated that the spectral features of LLMs’ attention maps, specifically the eigenvalues of the Laplacian matrix, carry a signal capable of detecting hallucinations. Specifically, we proposed the $\operatorname{LapEigvals}$ method, which employs the top-$k$ eigenvalues of the Laplacian as input to the hallucination detection probe. Through extensive evaluations, we empirically showed that our method consistently achieves state-of-the-art performance among all tested approaches. Furthermore, multiple ablation studies demonstrated that our method remains stable across varying numbers of eigenvalues, diverse prompts, and generation temperatures while offering reasonable generalisation.
In addition, we hypothesise that self-supervised learning (Balestriero et al., 2023) could yield a more robust and generalisable approach while uncovering non-trivial intrinsic features of attention maps. Notably, results such as those in Section 6.3 suggest intriguing connections to recent advancements in LLM research (Veličković et al., 2024; Barbero et al., 2024), highlighting promising directions for future investigation.
@@
Supervised method In our approach, one must provide labelled hallucinated and non-hallucinated examples to train the hallucination probe. While this can be handled by the llm-as-judge, it might introduce some noise or pose a risk of overfitting. Limited generalisation across LLM architectures The method is incompatible with LLMs having different head and layer configurations. Developing architecture-agnostic hallucination probes is left for future work. Minimum length requirement Computing $\operatorname{top-k}$ Laplacian eigenvalues demands attention maps of at least $k$ tokens (e.g., $k=100$ require 100 tokens). Open LLMs Our method requires access to the internal states of LLM thus it cannot be applied to closed LLMs. Risks Please note that the proposed method was tested on selected LLMs and English data, so applying it to untested domains and tasks carries a considerable risk without additional validation.
@@ Acknowledgements
We sincerely thank Piotr Bielak for his valuable review and insightful feedback, which helped improve this work. This work was funded by the European Union under the Horizon Europe grant OMINO – Overcoming Multilevel INformation Overload (grant number 101086321, https://ominoproject.eu/). Views and opinions expressed are those of the authors alone and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor the European Research Executive Agency can be held responsible for them. It was also co-financed with funds from the Polish Ministry of Education and Science under the programme entitled International Co-Financed Projects, grant no. 573977. This work was co-funded by the National Science Centre, Poland under CHIST-ERA Open & Re-usable Research Data & Software (grant number 2022/04/Y/ST6/00183).
@@ Appendix A Details of motivational study
We present a detailed description of the procedure used to obtain the results presented in Section 2 along with additional results for other datasets and LLM.
To test whether there is a statistically significant difference in the values of $\operatorname{AttnEigvals}$ and Laplacian eigenvalues, we first took QA datasets and ran inference with three LLMs, namely Llama-3.1-8BLlama-3.2-3B  and Phi-3.5 . Then, we extracted attention maps and computed $\operatorname{AttentionScore}$ (Sriramanan et al., 2024), i.e., log-determinant of attention maps. Unlike original work, we did not sum the scores over heads as we performed analysis at a single-head level of granularity. Also, we computed the Laplacian according to the definition presented in Section 3, took the 10 largest eigenvalues for each head, and treated each eigenvalue as a separate example. Finally, we ran the Mann–Whitney U test, leveraging SciPy implementation (Virtanen et al., 2020), and gathered $p$-values presented in Figure 1.
Table 3 presents the percentage of heads having a statistically significant difference in feature values between hallucinated and non-hallucinated examples, as indicated by $p<0.05$ from the Mann-Whitney U test. These results show that the Laplacian eigenvalues better distinguish between the two classes for all considered LLMs and datasets.
Table 3: Percentage of heads having a statistically significant difference in feature values between hallucinated and non-hallucinated examples, as indicated by $p<0.05$ from the Mann-Whitney U test. Results were obtained for $\operatorname{AttentionScore}$ and the 10 largest Laplacian eigenvalues on 6 datasets and 3 LLMs.
[…]
@@
Appendix B Details of QA datasets
In this work, we used 6 open and publicly available question answering datasets: NQ-Open (Kwiatkowski et al., 2019) (CC-BY-SA-3.0 license), SQuADv2 (Rajpurkar et al., 2018) (CC-BY-SA-4.0 license), TruthfulQA (Apache-2.0 license) (Lin et al., 2022), HALUEval (MIT license) (Li et al., 2023), CoQA (Reddy et al., 2019) (domain-dependent licensing, detailed on https://stanfordnlp.github.io/coqa/), while TriviaQA (lacks clear licensing information, but was primarily shared as public benchmark). Research purposes fall into the intended use of these datasets. To preprocess and filter TriviaQA, CoQA, and SQuADv2 we utilized open-source code of (Chen et al., 2024) [^5] https://github.com/alibaba/eigenscore (MIT license), which also borrows from (Farquhar et al., 2024) [^6] https://github.com/lorenzkuhn/semantic_uncertainty (MIT license). In Figure 8, we provide histogram plots of the number of tokens for $question$ and $answer$ of each dataset computed with meta-llama/Llama-3.1-8B-Instruct tokeniser.

### v2

@@ 6 Ablation studies
To better understand the behavior of our method under different conditions, we conduct a comprehensive ablation study. This analysis provides valuable insights into the factors driving the $\operatorname{LapEigvals}$ performance and highlights the robustness of our approach across various scenarios. In order to ensure reliable results, we perform all studies on the TriviaQA dataset, which has a moderate input size and number of examples.
@@ 6.1 How does the number of eigenvalues influence performance?
First, we verify how the number of eigenvalues influences the performance of the hallucination probe and present results for Mistral-Small-24B in Figure 4 (results for all models are showcased in Figure 10 in Appendix H). Generally, using more eigenvalues improves performance, but there is less variation in performance among different values of $k$ for $\operatorname{LapEigvals}$ compared to the baseline. Moreover, $\operatorname{LapEigvals}$ achieves significantly better performance with smaller input sizes, as $\operatorname{AttnEigvals}$ with the largest k=100k{=}100 fails to surpass $\operatorname{LapEigvals}$’s performance at k=5k{=}5. These results confirm that spectral features derived from the Laplacian carry a robust signal indicating the presence of hallucinations and highlight the strength of our method.
Figure 4: Probe performance across different top-$k$ eigenvalues: $k\in\{5,10,25,50,100\}$ for TriviaQA dataset with temp=1.0temp{=}1.0 and Mistral-Small-24B LLM.
@@ 6.2 Does using all layers at once improve performance?
Second, we demonstrate that using all layers of an LLM instead of a single one improves performance. In Figure 5, we compare per-layer to all-layer efficacy for Mistral-Small-24B (results for all models are showcased in Figure 11 in Appendix H). For the per-layer approach, better performance is generally achieved with deeper LLM layers. Notably, peak performance varies across LLMs, requiring an additional search for each new LLM. In contrast, the all-layer probes consistently outperform the best per-layer probes across all LLMs. This finding suggests that information indicating hallucinations is spread across many layers of LLM, and considering them in isolation limits detection accuracy. Further, Table 6 in Appendix G summarises outcomes for the two variants on all datasets and LLM configurations examined in this work.
Figure 5: Analysis of model performance across different layers for Mistral-Small-24B and TriviaQA dataset with temp=1.0temp{=}1.0 and k=100k{=}100 top eigenvalues (results for models operating on all layers provided for reference).
@@ 6.3 Does sampling temperature influence results?
Here, we compare $\operatorname{LapEigvals}$ to the baselines on hallucination datasets, where each dataset contains answers generated at a specific decoding temperature. Higher temperatures typically produce more hallucinated examples (Lee, 2023; Renze, 2024), leading to dataset imbalance. Thus, to mitigate the effect of data imbalance, we sample a subset of 1,0001{,}000 hallucinated and 1,0001{,}000 non-hallucinated examples $10$ times for each temperature and train hallucination probes. Interestingly, in Figure 6, we observe that all models improve their performance at higher temperatures, but $\operatorname{LapEigvals}$ consistently achieves the best accuracy on all considered temperature values. The correlation of efficacy with temperature may be attributed to differences in the characteristics of hallucinations at higher temperatures compared to lower ones (Renze, 2024). Also, hallucination detection might be facilitated at higher temperatures due to underlying properties of softmax function (Veličković et al., 2024), and further exploration of this direction is left for future work.
Figure 6: Test AUROC for different sampling temptemp values during answer decoding on the TriviaQA dataset, using k=100k{=}100 eigenvalues for $\operatorname{LapEigvals}$ and $\operatorname{AttnEigvals}$ with the Llama-3.1-8B LLM. Error bars indicate the standard deviation over 10 balanced samples containing $N=1000$ examples per class.
@@ 6.4 How does $\operatorname{LapEigvals}$ generalizes?
To check whether our method generalizes across datasets, we trained the hallucination probe on features from the training split of one QA dataset and evaluated it on the features from the test split of a different QA dataset. Due to space limitations, we present results for selected datasets and provide extended results and absolute efficacy values in Appendix I. Figure 7 showcases the percent drop in Test AUROC when using a different training dataset compared to training and testing on the same QA dataset. We can observe that $\operatorname{LapEigvals}$ provides a performance drop comparable to other baselines, and in several cases, it generalizes best. Interestingly, all methods exhibit poor generalization on TruthfulQA and GSM8K. We hypothesize that the weak performance on TruthfulQA arises from its limited size and class imbalance, whereas the difficulty on GSM8K likely reflects its distinct domain, which has been shown to hinder hallucination detection (Orgad et al., 2025). Additionally, in Appendix I, we show that $\operatorname{LapEigvals}$ achieves the highest test performance in all scenarios (except for TruthfulQA).
Figure 7: Generalization across datasets measured as a percent performance drop in Test AUROC (less is better) when trained on one dataset and tested on the other. Training datasets are indicated in the plot titles, while test datasets are shown on the $x$-axis. Results computed on Llama-3.1-8B with k=100k{=}100 top eigenvalues and temp=1.0temp{=}1.0. Results for all datasets are presented in Appendix I.
@@ 6.5 How does performance vary across prompts?
Lastly, to assess the stability of our method across different prompts used for answer generation, we compared the results of the hallucination probes trained on features regarding four distinct prompts, the content of which is included in Appendix M. As shown in Table 2, $\operatorname{LapEigvals}$ consistently outperforms all baselines across all four prompts. While we can observe variations in performance across prompts, $\operatorname{LapEigvals}$ demonstrates the lowest standard deviation ($0.05$) compared to $\operatorname{AttnLogDet}$ ($0.016$) and $\operatorname{AttnEigvals}$ ($0.07$), indicating its greater robustness.
Table 2: Test AUROC across four different prompts for answers on the TriviaQA dataset using Llama-3.1-8B  with temp=1.0temp{=}1.0 and k=50k{=}50 (some prompts have led to fewer than 100 tokens). Prompt $\boldsymbol{p_{3}}$ was the main one used to compare our method to baselines, as presented in Tables 1.
@@ 7 Related Work
White-box methods have emerged as a promising approach for detecting hallucinations (Farquhar et al., 2024; Azaria and Mitchell, 2023; Arteaga et al., 2024; Orgad et al., 2025). These methods are universal across all LLMs and do not require additional domain adaptation compared to black-box ones (Farquhar et al., 2024). They draw inspiration from seminal works on analyzing the internal states of simple neural networks (Alain and Bengio, 2016), which introduced linear classifier probes – models operating on the internal states of neural networks. Linear probes have been widely applied to the internal states of LLMs, notably for detecting hallucinations.
@@
One of the first such probes was SAPLMA (Azaria and Mitchell, 2023), which demonstrated that one could predict the correctness of generated text straight from LLM’s hidden states. Further, the INSIDE method (Chen et al., 2024) tackled hallucination detection by sampling multiple responses from an LLM and evaluating consistency between their hidden states using a normalized sum of the eigenvalues from their covariance matrix. Also, (Farquhar et al., 2024) proposed a complementary probabilistic approach, employing entropy to quantify the model’s intrinsic uncertainty. Their method involves generating multiple responses, clustering them by semantic similarity, and calculating Semantic Entropy using an appropriate estimator. To address concerns regarding the validity of LLM probes, (Marks and Tegmark, 2024) introduced a high-quality QA dataset with simple true/false answers and causally demonstrated that the truthfulness of such statements is linearly represented in LLMs, which supports the use of probes for short texts.
@@
Self-consistency methods (Liang et al., 2024), like INSIDE or Semantic Entropy, require multiple runs of an LLM for each input example, which substantially lowers their applicability. Motivated by this limitation, (Kossen et al., 2024) proposed to use Semantic Entropy Probe, which is a small model trained to predict expensive Semantic Entropy (Farquhar et al., 2024) from LLM’s hidden states. Notably, (Orgad et al., 2025) explored how LLMs encode information about truthfulness and hallucinations. First, they revealed that truthfulness is concentrated in specific tokens. Second, they found that probing classifiers on LLM representations do not generalize well across datasets, especially across datasets requiring different skills, which we confirmed in Section 6.4. Lastly, they showed that the probes could select the correct answer from multiple generated answers with reasonable accuracy, meaning LLMs make mistakes at the decoding stage, besides knowing the correct answer.
@@
Recent studies have started to explore hallucination detection exclusively from attention maps. (Chuang et al., 2024a) introduced the lookback ratio, which measures how much attention LLMs allocate to relevant input parts when answering questions based on the provided context. The work most closely related to ours is (Sriramanan et al., 2024), which introduces the $\operatorname{AttentionScore}$ method. Although the process is unsupervised and computationally efficient, the authors note that its performance can depend highly on the specific layer from which the score is extracted. Compared to $\operatorname{AttentionScore}$, our method is fully supervised and grounded in graph theory, as we interpret inference in LLM as a graph. While $\operatorname{AttentionScore}$ aggregates only the attention diagonal to compute its log-determinant, we instead derive features from the graph Laplacian, which captures all attention scores (see Eq. (1) and (2)). Additionally, we utilize all layers for detecting hallucination rather than a single one, demonstrating effectiveness of this approach. We also demonstrate that it performs poorly on the datasets we evaluated. Nonetheless, we drew inspiration from their approach, particularly using the lower triangular structure of matrices when constructing features for the hallucination probe.
@@ 8 Conclusions
In this work, we demonstrated that the spectral features of LLMs’ attention maps, specifically the eigenvalues of the Laplacian matrix, carry a signal capable of detecting hallucinations. Specifically, we proposed the $\operatorname{LapEigvals}$ method, which employs the top-$k$ eigenvalues of the Laplacian as input to the hallucination detection probe. Through extensive evaluations, we empirically showed that our method consistently achieves state-of-the-art performance among all tested approaches. Furthermore, multiple ablation studies demonstrated that our method remains stable across varying numbers of eigenvalues, diverse prompts, and generation temperatures while offering reasonable generalization.
In addition, we hypothesize that self-supervised learning (Balestriero et al., 2023) could yield a more robust and generalizable approach while uncovering non-trivial intrinsic features of attention maps. Notably, results such as those in Section 6.3 suggest intriguing connections to recent advancements in LLM research (Veličković et al., 2024; Barbero et al., 2024), highlighting promising directions for future investigation.
@@
Supervised method In our approach, one must provide labelled hallucinated and non-hallucinated examples to train the hallucination probe. While this can be handled by the llm-as-judge, it might introduce some noise or pose a risk of overfitting. Limited generalization across LLM architectures The method is incompatible with LLMs having different head and layer configurations. Developing architecture-agnostic hallucination probes is left for future work. Minimum length requirement Computing $\operatorname{top-k}$ Laplacian eigenvalues demands attention maps of at least $k$ tokens (e.g., k=100k{=}100 require 100 tokens). Open LLMs Our method requires access to the internal states of LLM thus it cannot be applied to closed LLMs. Risks Please note that the proposed method was tested on selected LLMs and English data, so applying it to untested domains and tasks carries a considerable risk without additional validation.
@@ Acknowledgements
We sincerely thank Piotr Bielak for his valuable review and insightful feedback, which helped improve this work. This work was funded by the European Union under the Horizon Europe grant OMINO – Overcoming Multilevel INformation Overload (grant number 101086321, https://ominoproject.eu/). Views and opinions expressed are those of the authors alone and do not necessarily reflect those of the European Union or the European Research Executive Agency. Neither the European Union nor the European Research Executive Agency can be held responsible for them. It was also co-financed with funds from the Polish Ministry of Education and Science under the programme entitled International Co-Financed Projects, grant no. 573977. We gratefully acknowledge the Wroclaw Centre for Networking and Supercomputing for providing the computational resources used in this work. This work was co-funded by the National Science Centre, Poland under CHIST-ERA Open & Re-usable Research Data & Software (grant number 2022/04/Y/ST6/00183). The authors used ChatGPT to improve the clarity and readability of the manuscript.
@@ Appendix A Details of motivational study
We present a detailed description of the procedure used to obtain the results presented in Section 2, along with additional results for other datasets and LLMs.
Our goal was to test whether $\operatorname{AttentionScore}$ and eigenvalues of Laplacian matrix (used by our $\operatorname{LapEigvals}$) differ significantly when examples are split into hallucinated and non-hallucinated groups. To this end, we used 7 datasets (Section 4.1) and ran inference with 5 LLMs (Section 4.1) using temp=0.1temp{=}0.1. From the extracted attention maps, we computed $\operatorname{AttentionScore}$ (Sriramanan et al., 2024), defined as the log-determinant of the attention matrices. Unlike the original work, we did not aggregate scores across heads, but instead analyzed them at the single-head level. For $\operatorname{LapEigvals}$, we constructed the Laplacian as defined in Section 3, extracted the 10 largest eigenvalues per head, and applied the same single-head analysis as for $\operatorname{AttnEigvals}$. Finally, we performed the Mann–Whitney U test (Mann and Whitney, 1947) using the SciPy implementation (Virtanen et al., 2020) and collected the resulting $p$-values
Table 3 presents the percentage of heads having a statistically significant difference in feature values between hallucinated and non-hallucinated examples, as indicated by $p<0.05$ from the Mann-Whitney U test. These results show that the Laplacian eigenvalues better distinguish between the two classes for almost all considered LLMs and datasets.
Table 3: Percentage of heads having a statistically significant difference in feature values between hallucinated and non-hallucinated examples, as indicated by $p<0.05$ from the Mann-Whitney U test. Results were obtained for $\operatorname{AttentionScore}$ and the 10 largest Laplacian eigenvalues on 6 datasets and 5 LLMs.
LLM Dataset % $\%\text{ of }p<0.05$
@@
Mistral-Nemo CoQA 35 78
Mistral-Nemo GSM8K 90 71
Mistral-Nemo HaluevalQA 78 82
Mistral-Nemo NQOpen 64 57
Mistral-Nemo SQuADv2 54 56
Mistral-Nemo TriviaQA 71 74
Mistral-Nemo TruthfulQA 40 50
Mistral-Small-24B CoQA 28 78
Mistral-Small-34B GSM8K 75 72
Mistral-Small-24B HaluevalQA 68 70
Mistral-Small-24B NQOpen 45 51
Mistral-Small-24B SQuADv2 75 82
Mistral-Small-24B TriviaQA 65 70
Mistral-Small-24B TruthfulQA 43 52
Appendix B Bounds of the Laplacian
In the following section, we prove that the Laplacian defined in 3 is bounded and has at least one zero eigenvalue. We denote eigenvalues as $\lambda_{i}$, and provide derivation for a single layer and head, which holds also after stacking them together into a single graph (set of per-layer graphs). For clarity, we omit the superscript (l,h){(l,h)} indicating layer and head.
Lemma 1.
The Laplacian eigenvalues are bounded: $1\leq\lambda_{i}\leq 1$.
Proof.
Due to the lower-triangular structure of the Laplacian, its eigenvalues lie on the diagonal and are given by:
$\lambda_{i}=\mathbf{L}_{ii}=d_{ii}-a_{ii}$
The out-degree is defined as:
$d_{ii}=\frac{\sum_{u}{a_{ui}}}{T-i},$
Since $0\leq a_{ui}\leq 1$, the sum in the numerator is upper bounded by TiT-i, therefore $d_{ii}\leq 1$, and consequently $\lambda_{i}=\mathbf{L}_{ii}\leq 1$, which concludes upper-bound part of the proof.
Recall that eigenvalues lie on the main diagonal of the Laplacian, hence $\lambda_{i}=\frac{\sum_{u}{a_{uj}}}{T-i}-a_{ii}$. To find the lower bound of $\lambda_{i}$, we need to minimize $X=\frac{\sum_{u}{a_{uj}}}{T-i}$ and maximize $Y=a_{ii}$. First, we note that $X$’s denominator is always positive Ti>0T-i>0, since $i\in\{0\dots(T-1)\}$ (as defined by Eq. (2)). For the numerator, we recall that $0\leq a_{ui}\leq 1$; therefore, the sum has its minimum at 0, hence $X\geq 0$. Second, to maximize $Y=a_{ii}$, we can take maximum of $0\leq a_{ii}\leq 1$ which is $1$. Finally, XY=1X-Y=-1, consequently $\mathbf{L}_{ii}\geq-1$, which concludes the lower-bound part of the proof. ∎
Lemma 2.
For every $\mathbf{L}_{ii}$, there exists at least one zero-eigenvalue, and it corresponds to the last token $T$, i.e., $\lambda_{T}=0$.
Proof.
Recall that eigenvalues lie on the main diagonal of the Laplacian, hence $\lambda_{i}=\frac{\sum_{u}{a_{uj}}}{T-i}-a_{ii}$. Consider last token, wherein the sum in the numerator reduces to $\sum_{u}{a_{uj}}=a_{TT}$, denominator becomes Ti=T(T1)=1T-i=T-(T-1)=1, thus $\lambda_{T}=\frac{a_{TT}}{1}-a_{TT}=0$. ∎
Appendix C Implementation details
In our experiments, we used HuggingFace Transformers (Wolf et al., 2020), PyTorch (Ansel et al., 2024), and scikit-learn (Pedregosa et al., 2011). We utilized Pandas (team, 2020) and Seaborn (Waskom, 2021) for visualizations and analysis. To version data, we employed DVC (Kuprieiev et al., 2025). The Cursor IDE was employed to assist with code development. We performed LLM inference and acquired attention maps using a single Nvidia A40 with 40GB VRAM, except for Mistral-Small-24B for which we used Nvidia H100 with 96GB VRAM. Training the hallucination probe was done using the CPU only. To compute labels using the llm-as-judge approach, we leveraged gpt-4o-mini model available through OpenAI API. Detailed hyperparameter settings and code to reproduce the experiments are available in the public Github repository: https://github.com/graphml-lab-pwr/lapeigvals.
Appendix D Details of QA datasets
We used 7 open and publicly available question answering datasets: NQ-Open (Kwiatkowski et al., 2019) (CC-BY-SA-3.0 license), SQuADv2 (Rajpurkar et al., 2018) (CC-BY-SA-4.0 license), TruthfulQA (Apache-2.0 license) (Lin et al., 2022), HaluEvalQA (MIT license) (Li et al., 2023), CoQA (Reddy et al., 2019) (domain-dependent licensing, detailed on https://stanfordnlp.github.io/coqa/), TriviaQA (Apache 2.0 license), GSM8K (Cobbe et al., 2021)(MIT license). Research purposes fall into the intended use of these datasets. To preprocess and filter TriviaQA, CoQA, and SQuADv2 we utilized open-source code of (Chen et al., 2024) [^7] https://github.com/alibaba/eigenscore (MIT license), which also borrows from (Farquhar et al., 2024) [^8] https://github.com/lorenzkuhn/semantic_uncertainty (MIT license). In Figure 8, we provide histogram plots of the number of tokens for questionquestion and answeranswer of each dataset computed with meta-llama/Llama-3.1-8B-Instruct tokenizer.